# Resolving UnderEdit & OverEdit with Iterative & Neighbor-Assisted Model Editing

### v1

@@ Abstract
Large Language Models (LLMs) are used in various downstream language tasks, making it crucial to keep their knowledge up-to-date, but both retraining and fine-tuning the model can be costly. Model editing offers an efficient and effective alternative by a single update to only a key subset of model parameters. While being efficient, these methods are not perfect. Sometimes knowledge edits are unsuccessful, i.e., UnderEdit, or the edit contaminated neighboring knowledge that should remain unchanged, i.e., OverEdit. To address these limitations, we propose iterative model editing, based on our hypothesis that a single parameter update is often insufficient, to mitigate UnderEdit, and neighbor-assisted model editing, which incorporates neighboring knowledge during editing to minimize OverEdit. Extensive experiments demonstrate that our methods effectively reduce UnderEdit up to 38 percentage points and OverEdit up to 6 percentage points across multiple model editing algorithms, LLMs, and benchmark datasets.
@@
Iterative & Neighbor-Assisted Model Editing
Bhiman Kumar Baghel  Scott M. Jordan  Zheyuan Ryan Shi  Xiang Lorraine Li Department of Computer Science, University of Pittsburgh, PA, USA {bkb45, xianglli}@pitt.edu
@@ 1 Introduction
LLMs have been widely used as repositories of factual and specialized knowledge (Petroni et al., 2020; Jiang et al., 2021; Roberts et al., 2020; Youssef et al., 2023). However, the world is constantly changing, with knowledge and information evolving rapidly, such as significant government policy changes and their wide impacts across various domains. Thus, it is essential for many NLP applications, such as text generation, question answering, and knowledge retrieval, to have models that can adapt to knowledge changes both effectively and efficiently. Re-training an LLM is resource-intensive (Patterson et al., 2021). Standard supervised fine-tuning is data hungry and less effective (Meng et al., 2023b). Model-editing, which directly modifies important model parameters for making the prediction, has emerged as a more efficient alternative for updating outdated information (Meng et al., 2023a, b; Li et al., 2024). These methods adopt a “locate and edit” approach, wherein they first identify the parameter locations associated with outdated knowledge and then update the parameters to enable the model to incorporate and predict the new knowledge.
@@
In summary, we propose solutions to two fundamental challenges in model editing: UnderEdit, where edits fail, and OverEdit, where neighboring knowledge is erroneously modified. We evaluate our approach using three “locate and edit” model editing algorithms, ROME (Meng et al., 2023a), MEMIT (Meng et al., 2023b), and PMET (Li et al., 2024), and applied to three LLMs: GPT-2 XL (1.5B) (Radford et al., 2019), GPT-J (6B) (Wang and Komatsuzaki, 2021), and Llama 2 (7B) (Touvron et al., 2023). Our experiments are conducted on two widely used factual knowledge editing benchmarks: COUNTERFACT (Meng et al., 2023a) and ZsRE (Levy et al., 2017). Our results show that iterative model editing improves edit success while also reducing the approximation error introduced by the rank-one update. Furthermore, we demonstrate that incorporating neighboring knowledge during model editing leads to fewer unintended modifications to neighboring knowledge at test time, resulting in stronger edit performance.
2 Background
In this section, we provide background on the locate and edit model editing framework along with the notation used throughout the paper.
@@ 2 Background
The weight update is performed in two stages: Optimization stage finds ideal values for the network’s hidden state in certain transformer layer to make $o^{*}$ likely and Spread stage updates the weights of the last MLP in casual layer(s) to approximate this ideal hidden state. We detail these stages below for MEMIT (Meng et al., 2023b) and discuss its differences to PMET (Li et al., 2024) and ROME (Meng et al., 2023a).
@@
The goal of the Optimization stage is to find what outputs in the causal layers would lead to a high likelihood on $o^{*}$. The methods we investigate search for ideal outputs in different locations. MEMIT searches for an ideal output, $\bar{h}^{l_{c}}_{t}$, for the last casual layer at $t$ the last token index of the subject $s$. The search is performed by finding a vector $\delta$ to add to current hidden state value $h^{l_{c}}_{t}$. We represent the output probability distribution of the model using $h^{l_{c}}_{t}$ and $\delta$ as $f_{\theta}(x,h^{l_{c}}_{t}+\delta)$.
@@
The ideal hidden state for the prompt $x=(s,r)$ is $\bar{h}_{t}^{l_{c}}=h_{t}^{l_{c}}+\delta^{*}$, where $\delta^{*}$ is found by performing gradient descent on $g$. This ideal hidden state is then used in computing weight update in the next stage. PMET differs from MEMIT by searching for an ideal outputs for the attention module and MLP modules in layer $l_{c}$. ROME searches for an ideal output for the MLP module of a single layer in the set of causal layers.
Algo Optimization Spread
MEMIT $h^{l_{c}}$ $W^{l_{1}}\cdots W^{l_{c}}$
PMET $\textit{Attn}^{l_{c}}$ $\textit{z}^{l_{c}}$ $W^{l_{1}}\cdots W^{l_{c}}$
ROME $z^{l_{*}}$ $W^{l_{*}}$
Figure 2: The table compares three memory editing algorithms, each of which optimizes different parameters within a single transformer layer and propagates changes across causal layer(s). Despite their differences, all algorithms follow the same Optimization and Spread stages. The figure below shows a simplified transformer layer, composed of an attention module and MLP modules. Only the last MLP module is shown, as all algorithms modify parameters within it.
@@
The goal of the Spread stage is to find new weights $\theta^{\prime}$ such that the hidden state after the update $\hat{h}^{l_{c}}_{t}$is close to the ideal hidden state $\bar{h}^{l_{c}}_{t}$ for all desired edits in $D$. Not all weights in the network are updated, only the weights $W^{l}$ corresponding to the weights of the last MLP layer in causal layers are updated. The weight update methods are derived from a rank-one approximation to make $\hat{h}^{l_{c}}_{t}\approx\bar{h}^{l_{c}}_{t}$, which can lead to failure. The different algorithms also update a different set of weights; MEMIT and PMET update $W^{l}$ for all causal layers, while ROME only updates one $W^{l}$. The algorithm differences are summarized in Figure 2.
@@ 3.1 Iterative Model Editing
There are two possible reasons for UnderEdit to occur. The first is that $\bar{h}^{l_{c}}_{t}$ does not reflect a hidden state for a successful edit. The second is that the weight update results in $\hat{h}^{l_{c}}_{t}\not\approx\bar{h}^{l_{c}}_{t}$. We hypothesize that both of these potential problems can be addressed by running the memory edit process multiple times because: 1) it allows for potentially finding better $\bar{h}^{l_{c}}_{t}$ after updating the model parameters so that $\|\hat{h}_{t}^{l_{c}}-\bar{h}^{l_{c}}_{t}\|\leq\|h^{l_{c}}_{t}-\bar{h}^{l_{c}}_{t}\|$, and 2) on the next iteration, the approximation used in the Spread stage for the weight update will be better since $\hat{h}^{t}_{l_{c}}$ is closer to $\bar{h}_{t}^{l_{c}}$ than $h_{t}^{l_{c}}$. We detail this iterative process below for MEMIT, but it can also be adapted to ROME and PMET by replacing $\bar{h}_{t}^{l_{c}}$ with the targets of the optimization procedure for those algorithms.
@@
The iterations end when the model perplexity using $\hat{h}_{t,k}^{l_{c}}$ is within $\epsilon$ of the perplexity using $\bar{h}_{t,k}^{l_{c}}$, i.e.,
@@
Figure 3 illustrates how iterative model editing allows $\hat{h}^{l_{c}}_{t,k}$ to have a perplexity closer to $\bar{h}_{t,k}^{l_{c}}$ over time. The figure also makes it clear that most of the improvement of iterative model editing comes from applying Spread multiple times since the perplexity of $\bar{h}_{t,k}^{l_{c}}$ does not change as much.
@@ 3.2 Neighbor-Assisted Model Editing
Model editing not only needs to modify the model’s output from $o$ to $o^{*}$ given $(s,r)$, but also preserve the model’s output for neighboring knowledge, i.e $(\tilde{s},r,o)$, where $\tilde{s}$ is a new subject sharing the same relation $r$. An example of this preservation is shown in Figure 1a: iPhone 11 [$\tilde{s}$] is still produced by [$r$] Apple [$o$] despite iPad is edited to produced by Honda [^1] In the COUNTERFACT dataset. Existing model editing algorithms struggle to preserve neighboring knowledge because Optimization is designed solely to maximize the likelihood of the new knowledge, $(s,r,o^{*})$. Moreover, iterative model editing can exacerbate this OverEdit issue, as each iteration continues to reinforce the new knowledge without explicitly preserving neighboring knowledge.
@@
Table 1: Iterative model editing results on COUNTERFACT for at most 10 iterations (denoted by k). We compare the evaluation metrics of iteration that met stopping criterion $|\Delta p_{k}|\leq 1$ to that of their corresponding first iteration and bold the higher value. PMET on GPT-2 XL require more than 5 iterations to achieve our stopping criteria. Results for all iterations are provided in Table 6 (Appendix D). While ROME is known to collapse (results reported in Table 6), we observed a unique case of collapse with Llama-2 (7B) specifically when using MEMIT. We discuss this in Section 5.
@@ 4 Experimental Details
In this section, we detail the experiments to demonstrate the effectiveness of iterative and neighbor-assisted model editing with MEMIT, PMET, and ROME. We evaluate these algorithms with our modifications across three LLMs: GPT-2 XL (1.5B), GPT-J (6B), and Llama-2 (7B). We use the EasyEdit [^2] https://github.com/zjunlp/EasyEdit framework Wang et al. (2024b) with their default hyperparameters. Additional implementation details are provided in Appendix B.
@@ 5.1 Iterative Model Editing Results
First, iterative model editing consistently improves performance, with the overall success scores increasing across iterations for most models and algorithms. The overall success improvement stems from enhanced efficacy and generalization capabilities, which means fewer cases of UnderEdit. Specifically, we observed an increase in success accuracy of up to 38 percentage points, with a greater improvement in efficacy accuracy of up to 77 percentage points (PMET on GPT-2 XL). We conducted more analysis in section 6.1 to showcase the efficacy improvement is mostly coming from UnderEdit examples. Secondly, as iteration goes, the perplexity difference constantly goes down in most cases. Finally, the proposed stopping criterion ($|\Delta p_{k}|\leq\epsilon$) consistently halts the process, validating its reliability. Using this criterion also yields better overall scores compared to executing the algorithm only once.
@@
Although iterative editing is effective in most cases, we also observed model collapse as editing progresses, indicated by high model perplexity on ME-PPL-50. This collapse behavior aligns with the continuous editing failures observed in previous work Gupta et al. (2024); Meng et al. (2023a). This suggests that when the combination of the model and editing algorithm succeeds in continuous editing, an essential experimental setting, iterative editing can further improve model performance. Specifically, ROME collapse under iterative 333 ROME does not support batch editing, as it can only modify one fact at a time Meng et al. (2023b). We discuss this further in Appendix A. editing, which is consistent with prior work’s finding on its inability to do continuous editing Gupta et al. (2024). We also found that Llama-2 (7B) collapse only when edited with MEMIT, similar finding fromYang et al. (2024) confirmed this. We hypothesize two contributing factors: (1) model-specific characteristics, as no GPT models exhibited similar collapse behavior, suggesting that certain training-related strategy may be more vulnerable to instability, and (2) MEMIT’s less precise weight updates compared to PMET, which may have introduced excessive parameter changes (Appendix A).

### v2

@@ Abstract
Large Language Models (LLMs) are widely deployed in downstream tasks, but keeping their knowledge up-to-date via retraining or fine-tuning is often computationally expensive. Model editing provides a more efficient alternative by updating a targeted subset of parameters, which often follows the locate-and-edit paradigm. Despite this efficiency, existing methods are limited: edits may fail to inject knowledge (UnderEdit) or unintentionally disrupt unrelated neighboring knowledge (OverEdit). To address these challenges, we propose two complementary methods: iterative model editing, which applies successive edits to mitigate UnderEdit, and neighbor-assisted model editing, which incorporates neighboring knowledge during editing to reduce OverEdit. Our extensive experiments show that these techniques improve editing performance across multiple LLMs, algorithms, and benchmarks, reducing UnderEdit by up to 38 percentage points and OverEdit by up to 6, while remaining broadly applicable to any locate-and-edit method.
@@
Iterative & Neighbor-Assisted Model Editingthanks: Preprint under-review
Bhiman Kumar Baghel  Scott M. Jordan  Zheyuan Ryan Shi  Xiang Lorraine Li Department of Computer Science, University of Pittsburgh, PA, USA {bkb45, scott.jordan, ryanshi, xianglli}@pitt.edu
@@ 1 Introduction
LLMs have been widely used as repositories of factual and specialized knowledge (Petroni et al., 2020; Jiang et al., 2021; Roberts et al., 2020; Youssef et al., 2023). However, the world is constantly changing, with knowledge and information evolving rapidly, such as significant government policy changes and their wide impacts across various domains. Thus, it is essential for many NLP applications, such as text generation, question answering, and knowledge retrieval, to have models that can adapt to knowledge changes both effectively and efficiently. Re-training an LLM is resource-intensive (Patterson et al., 2021). Standard supervised fine-tuning is data hungry and less effective (Meng et al., 2023b). Model-editing, which directly modifies important model parameters for making the prediction, has emerged as a more efficient alternative for updating outdated information (Meng et al., 2023a, b; Li et al., 2024; Fang et al., 2025). These methods adopt a “locate-and-edit” approach, where they first identify the parameter locations associated with outdated knowledge and then update the parameters to enable the model to incorporate and predict the new knowledge.
@@
In summary, we propose solutions to two fundamental challenges in model editing: UnderEdit, where edits fail, and OverEdit, where neighboring knowledge is erroneously modified. We evaluate our approach using four “locate and edit” model editing algorithms, ROME (Meng et al., 2023a), MEMIT (Meng et al., 2023b), PMET (Li et al., 2024) and AlphaEdit (Fang et al., 2025), and applied to four LLMs: GPT-2 XL (1.5B) (Radford et al., 2019), GPT-J (6B) (Wang and Komatsuzaki, 2021), Llama-2 (7B) (Touvron et al., 2023), and Llama-3.1 (8B) (Meta, 2024). Our experiments are conducted on two widely used factual knowledge editing benchmarks: COUNTERFACT (Meng et al., 2023a) and ZsRE (Levy et al., 2017). Our results show that iterative model editing improves edit success while also reducing the approximation error introduced by the rank-one update. Furthermore, we demonstrate that incorporating even a single neighboring knowledge during model editing reduces unintended modifications to neighboring knowledge at test time, resulting in stronger edit performance. Overall, our proposed methods are broadly applicable and consistently effective across current locate-and-edit approaches that adopt the two-stage editing framework, and are readily applicable to future methods built on this foundation.
2 The Locate-and-Edit Framework
In this section, we provide background on the locate-and-edit model editing framework along with the notation used throughout the paper.
@@ 2 The Locate-and-Edit Framework
The weight update is performed in two stages: Optimization stage finds ideal values for the network’s hidden state in certain transformer layer to make $o^{*}$ likely and Spread stage updates the weights of the last MLP in casual layer(s) to approximate this ideal hidden state. We detail these stages below for MEMIT and discuss its differences to PMET, AlphaEdit, and ROME.
Algorithm Optimization Stage Spread Stage
MEMIT (Meng et al., 2023b) $h^{l_{c}}$ (hidden state) $W^{l_{1}}\cdots W^{l_{c}}$ (least-squares update)
PMET (Li et al., 2024) $\textit{Attn}^{l_{c}}$ $\textit{z}^{l_{c}}$ (ideal attention + MLP) $W^{l_{1}}\cdots W^{l_{c}}$ (attention-free update)
AlphaEdit (Fang et al., 2025) $h^{l_{c}}$ (projected to null space) $W^{l_{1}}\cdots W^{l_{c}}$ (null space-constrained update)
ROME (Meng et al., 2023a) $z^{l_{*}}$ (ideal value) $W^{l_{*}}$ (rank-one equality-constrained update)
R-ROME (Gupta et al., 2024a) $z^{l_{*}}$, $k_{*}$ (averaged over context) $W^{l_{*}}$ (stabilized rank-one update)
EMMET (Gupta et al., 2024c) $h^{l_{c}}$ (batched hidden states) $W^{l_{1}}\cdots W^{l_{c}}$ (equality-constrained batch update)
ENCORE (Gupta et al., 2025) $h^{l_{c}}$ (with MPES early stopping) $W^{l_{1}}\cdots W^{l_{c}}$ (norm-constrained update)
EVOKE (LTI) Zhang et al. (2025) $h^{l_{c}}$ (constrained via multi-stage loss) Follows base method (e.g., ROME or MEMIT)
Table 1: We present a unifying overview of existing locate-and-edit algorithms within the two-stage framework of Optimization and Spread. This abstraction allows our proposed methods—iterative and neighbor-assisted editing—to be applied broadly across all listed algorithms, as well as to future methods built upon this framework. For detailed descriptions of each algorithm, see Appendix A.
@@
The goal of the Optimization stage is to find what outputs in the causal layers would lead to a high likelihood on $o^{*}$. The methods we investigate search for ideal outputs in different locations. MEMIT searches for an ideal output, $\bar{h}^{l_{c}}_{t}$, for the last casual layer at $t$, the last token index of the subject $s$. The search is performed by finding a vector $\delta$ to add to current hidden state value $h^{l_{c}}_{t}$. We represent the output probability distribution of the model using $h^{l_{c}}_{t}$ and $\delta$ as $f_{\theta}(x,h^{l_{c}}_{t}+\delta)$.
@@
The ideal hidden state for the prompt $x=(s,r)$ is $\bar{h}_{t}^{l_{c}}=h_{t}^{l_{c}}+\delta^{*}$, where $\delta^{*}$ is found by performing gradient descent on $g$. This ideal hidden state is then used in computing weight update in the next stage. AlphaEdit is the same as MEMIT, whereas PMET differs from MEMIT by searching for ideal outputs for the attention module and MLP modules in layer $l_{c}$. ROME searches for an ideal output for the MLP module of a single layer in the set of causal layers.
Figure 2: The diagram shows a simplified transformer layer to complement Table 1, composed of attention and MLP modules. Only the last MLP is shown, as all methods modify its parameters.
@@
The goal of the Spread stage is to find new weights $\theta^{\prime}$ such that the hidden state after the update $\hat{h}^{l_{c}}_{t}$ is close to the ideal hidden state $\bar{h}^{l_{c}}_{t}$ for all desired edits in $D$. Not all weights in the network are updated, only the weights $W^{l}$ corresponding to the weights of the last MLP layer in causal layers are updated. The weight update methods are derived from a rank-one approximation to make $\hat{h}^{l_{c}}_{t}\approx\bar{h}^{l_{c}}_{t}$, which can lead to failure. The different algorithms update different set of weights: MEMIT, PMET, and AlphaEdit update $W^{l}$ for all causal layers, whereas ROME only updates one $W^{l}$. Among these, AlphaEdit differs by computing $\theta^{\prime}$ using the null space projection method (Wang et al., 2021). The method differences are shown in Table 1, with Figure 2 complementing it.
Several other locate-and-edit methods exist, such as R-ROME (Gupta et al., 2024a), EMMET (Gupta et al., 2024c), EVOKE (LTI) (Zhang et al., 2025), and ENCORE (Gupta et al., 2025), all of which follow the same two-stage procedure (detailed in Appendix A). Below, we discuss the fundamental limitations of this framework and propose a general approach that applies broadly to all methods in this category—not just the four highlighted here, as summarized in Table 1.
@@ 3.1 Iterative Model Editing
There are two possible reasons for UnderEdit to occur. The first is that $\bar{h}^{l_{c}}_{t}$ does not reflect a hidden state for a successful edit. The second is that the weight update results in $\hat{h}^{l_{c}}_{t}\not\approx\bar{h}^{l_{c}}_{t}$. We hypothesize that both of these potential problems can be addressed by running the memory edit process multiple times because: 1) it allows for potentially finding better $\bar{h}^{l_{c}}_{t}$ after updating the model parameters so that $\|\hat{h}_{t}^{l_{c}}-\bar{h}^{l_{c}}_{t}\|\leq\|h^{l_{c}}_{t}-\bar{h}^{l_{c}}_{t}\|$, and 2) on the next iteration, the approximation used in the Spread stage for the weight update will be better since $\hat{h}^{t}_{l_{c}}$ is closer to $\bar{h}_{t}^{l_{c}}$ than $h_{t}^{l_{c}}$. We detail this iterative process below for MEMIT, but it can also be adapted to ROME, PMET and AlphaEdit by replacing $\bar{h}_{t}^{l_{c}}$ with the targets of the optimization procedure for those algorithms.
@@
The iterations end when model perplexity using $\hat{h}_{t,k}^{l_{c}}$ is within $\epsilon$ of the perplexity using $\bar{h}_{t,k}^{l_{c}}$, i.e.,
@@
Figure 3 illustrates how iterative model editing progressively brings $\hat{h}^{l_{c}}_{t,k}$ closer in perplexity to $\bar{h}_{t,k}^{l_{c}}$. The figure also highlights that most of the improvement stems from applying Spread multiple times, as the perplexity of $\bar{h}_{t,k}^{l_{c}}$ changes relatively little across iterations. However, the stability in perplexity does not imply that $\bar{h}_{t,k}^{l_{c}}$ remains identical at each iteration. We verified this through an additional experiment using iterative Spread, where Optimization was performed only in the first iteration. We found that keeping $\bar{h}_{t,k}^{l_{c}}$ fixed across iterations leads to overfitting and performance degradation, underscoring the necessity of running Optimization at every step. This confirms that, despite similar perplexities, the hidden states evolve across iterations and must be recomputed to ensure stable and effective edits. A detailed explanation of why the hidden states evolve across iterations is provided in Appendix D.
@@ 3.2 Neighbor-Assisted Model Editing
Model editing must not only change the model’s output from $o$ to $o^{*}$ given $(s,r)$, but also preserve outputs for neighboring knowledge, i.e $(\tilde{s},r,o)$, where $\tilde{s}$ is a new subject sharing the same relation $r$. A preservation example is shown in Figure 1a: iPhone 11 [$\tilde{s}$] is still produced by [$r$] Apple [$o$] despite iPad is edited to produced by Honda [^1] In the COUNTERFACT dataset.
Existing model editing algorithms struggle to preserve neighboring knowledge because Optimization is designed solely to maximize the likelihood of the new knowledge, $(s,r,o^{*})$. Moreover, iterative model editing can exacerbate this OverEdit issue, as each iteration continues to reinforce the new knowledge without explicitly preserving neighboring knowledge.
@@
AlphaEdit 1 94.33% 97.00% 47.67% 67.33% 59.00% 80.00% 61.67% 79.67% 30.79 18.84
2 100.00% 100.00% 67.67% 89.33% 51.67% 77.33% 68.00% 87.67% 31.47 0.15
Llama-3.1 (8B) Unedited 0 1.00% 7.00% 1.00% 9.33% 89.67% 11.33% 71.73
ROME 1 1.00% 78.00% 0.00% 68.00% 0.00% 66.00% 1.00% 70.00% 1.03E+05 1.44E+08
MEMIT 1 96.33% 98.00% 52.67% 80.33% 81.00% 98.00% 72.33% 91.33% 72.09 4,550.10
3 100.00% 100.00% 68.67% 93.67% 74.67% 97.00% 79.00% 96.67% 72.14 0.01
PMET 1 2.00% 96.00% 7.00% 78.00% 69.00% 98.00% 4.00% 90.00% 77.06 2,600.30
3 1.00% 98.00% 10.00% 88.00% 64.00% 97.00% 3.00% 94.00% 79.95 0.06
AlphaEdit 1 94.67% 97.00% 50.33% 78.33% 76.00% 97.00% 69.00% 90.00% 71.81 4,470.02
3 100.00% 100.00% 68.00% 93.67% 67.33% 95.67% 76.00% 96.67% 72.30 0.01
Table 2: Iterative model editing results on COUNTERFACT for at most 10 iterations (denoted by k). We compare the evaluation metrics of iteration that met stopping criterion $|\Delta p_{k}|\leq 1$ to that of their corresponding first iteration and bold the higher value. PMET on GPT-2 XL require more than 5 iterations to achieve our stopping criteria. Results for all iterations are provided in Table 9 (Appendix G). While ROME is known to collapse (results reported in Table 9), we observed a unique case of collapse with Llama-2 (7B) specifically when using MEMIT. We discuss this in Section 5. Note: Results for PMET on LLaMA-3.1 are based on a single subset, as results for the remaining subsets are still being generated. Results for AlphaEdit on GPT-J are currently unavailable due to an implementation error in EasyEdit, which we are actively working to resolve.
@@ 4 Experimental Details
In this section, we detail the experiments to demonstrate the effectiveness of iterative and neighbor-assisted model editing with MEMIT, PMET, AlphaEdit and ROME. We evaluate these algorithms with our modifications across four LLMs: GPT-2 XL (1.5B), GPT-J (6B), Llama-2 (7B) and Llama-3.1 (8B). We use EasyEdit [^2] https://github.com/zjunlp/EasyEditWang et al. (2024b) with default hyperparameters; implementation details are in AppendixB.
@@ 5.1 Iterative Model Editing Results
First, iterative model editing consistently improves performance, with the overall success scores increasing across iterations for most models and algorithms. The overall success improvement stems from enhanced efficacy and generalization capabilities, which means fewer cases of UnderEdit. Specifically, we observed an increase in success accuracy of up to 38 percentage points, with a greater improvement in efficacy accuracy of up to 77 percentage points (PMET on GPT-2 XL). We conducted more analysis in Appendix F.1 to showcase the efficacy improvement is mostly coming from UnderEdit examples. Secondly, as iteration goes, the perplexity difference constantly goes down in most cases. Finally, the proposed stopping criterion ($|\Delta p_{k}|\leq\epsilon$) consistently halts the process, validating its reliability. Using this criterion also yields better overall scores compared to executing the algorithm only once. Additionally, in Section 5.3, we compare our stopping criterion against two alternatives and find it to be the most effective.
@@
Although iterative editing is effective in most cases, we also observed model collapse as editing progresses, indicated by high model perplexity on ME-PPL-50 (5 out of 15 experiment settings). This collapse behavior aligns with the continuous editing failures observed in previous work Gupta et al. (2024b); Meng et al. (2023a). This suggests that when the combination of the model and editing algorithm succeeds in continuous editing, an essential experimental setting, iterative editing can further improve model performance. Specifically, ROME collapses under iterative 333 ROME does not support batch editing, as it can only modify one fact at a time Meng et al. (2023b). We discuss this further in Appendix A. editing, which aligns with prior findings on its inability to support continuous editing Gupta et al. (2024b). We also found that Llama-2 (7B) collapses only when edited with MEMIT, a result consistent with findings in Yang et al. (2024). Thus, we conclude that iterative model editing does not inherently lead to collapse. However, unstable models that fail with sequential editing may not benefit from this approach. Although investigating collapse is not the primary focus of this work, our findings (detailed in Appendix E) offer a useful foundation for future research in this area.
We also consistently observed a trade-off between generalization and specificity in Table 2. To further investigate this phenomenon, we focused on AlphaEdit, which aims to address this trade-off using null space projection to constrain model updates so that they minimally interfere with existing knowledge. Our results show that although AlphaEdit is designed to mitigate OverEdit, its performance still degrades in specificity under iterative model editing—indicating that this method does not fully resolve the problem. This persistent trade-off affects all locate-and-edit algorithms, and the proposed iterative editing is effective and widely applicable across this family of methods.